# Non-Iterative Disentangled Unitary Coupled-Cluster based on Lie-algebraic structure

### v2

@@ II Computational Details
To generate the symmetry-preserving MCP excitations, which is the final step of the fourth component needed for the completion of the NI-DUCC algorithm, we used an additional python code, that generates an MCP by ensuring that the pool of excitations meets the required completeness conditions. This involves computing the product group and building their corresponding Lie algebra (see the python code given in [60]). It is important to note that computing the product group of a given pool is resource-intensive, especially as the number of qubits increases. Consequently, Section IV.2 provides detailed information about the algorithm computational resource demands while increasing the number of qubits. Finally, to address this issue in order to better estimate the computational requirements, we converted completely the python code into a more efficient C++ capable of a faster MCP generation.
@@ III.1 Numerical Simulations
[…]
Figure 2: Energy convergence plots for ground states of H6 (12 qubits) and BeH2 (14 qubits), using the STO-3G basis set, at bond distances ( $r_{\text{H-H}}$ = 1.0 Å,) and ($r_{\text{Be-H}}$ = 3.5 Å), respectively. The plots are obtained using the NI-DUCC-VQE algorithm: Top panel (a, b), compares the quality of excitation set, insuring the significance of strong symmetric excitations when they are combined with the closure Lie algebraic properties. Panel (c, d), illustrates the rapid convergence as the number of layers, $k$, increases in the circuit. Bottom Panel (e, f), shows the fidelity property, which is calculated by overlapping $\langle\Psi_{j}(\vec{\theta^{*}})|\Psi_{g}\rangle$ between the computed NI-DUCC state at each optimization step ``$j^{\prime\prime}$ and the theoretical ground state $|\Psi_{g}\rangle$ of Hamiltonian ($\hat{H}$), where $|\Psi_{g}\rangle$ is obtained after computing the matrix diagonalization of $\hat{H}$.
@@
In the last protocol, where strongly symmetric excitations are considered as starters and incorporate Lie closure relations, we observe that NI-DUCC-VQE allows for a “full” convergence of the H6 and BeH2 ground states, with an error of less than 10-10 Hartree relative to the reference FCI energy. The convergence is very rapid, requiring only about 800 function evaluations. These observations suggest that strong symmetric excitations with closure algebraic relations, might approximate strongly correlated states, much better than the NI-DUCC ansatz formed from strongly symmetric excitations missing Lie algebraic closure relations. Moreover, since we simulated BeH2 at a larger bond length of 3.5 Å, where the ground state is more strongly correlated, we expected to see some difference in the convergence behavior with respect to H6, which was tested near equilibrium at r= 1.0 Å. However, this was not the case, indicating that the Lie algebraic closure relations help to overcome this challenge. This demonstrates that incorporating Lie algebraic closure with strong symmetric starters in the NI-DUCC wavefunction systematically provides a solution to the ordering problem. Indeed, several studies [29, 66, 67] demonstrate that addressing ordering problems can enhance convergence and improve UCC-VQE energy. Similarly, we have shown improvements in the NI-DUCC accuracy thanks to the use of Lie algebraic closure relations. We then further investigated the importance of using the “fourth” protocol in constructing the NI-DUCC wavefunction at $k=8$. As depicted in Figure 2 for H6 and Figure 2 for BeH2, we included energy convergence plots not only for $k=8$, but also for $k=3$ to $k=7$. We observed that when $k=3$ and $k=4$ are used, NI-DUCC approaches chemical accuracy (10-3 Hartree), while when using $k=5$, $k=6$ and $k=7$, it achieves an accuracy about $10^{-5}$ Hartree. Although with $k=8$, NI-DUCC-VQE approaches exact FCI solution (with an error less than 10-11 Hartree) after approximately 900 function evaluations without revealing any long plateaus, the differences between various $k$ layers may only matter, if one stops the calculation before a full convergence is reached. This could be the case, for instance, if achieving chemical accuracy is sufficient for a particular molecule. However, if the aim of the computation is to reach full convergence with the highest level of accuracy, then, using NI-DUCC excitations with sufficient number of layers is advantageous. Additionally, we have plotted (see Figure. 2 for H6 and Figure. 2 for BeH2), the overlap (“fidelity”) of the NI-DUCC-VQE state at each optimization step with the full configuration interaction (FCI) state, against the number of function evaluations (optimization steps). Various NI-DUCC tests were performed using different setups: (i) when strong symmetric starters are within closure Lie algebraic relations, we computed NI-DUCC($k=4$), NI-DUCC($k=6$), NI-DUCC($k=8$); (ii) when weak symmetric starters are within closure Lie algebraic relations, we computed NI-DUCC($k=8$). We observed that the NI-DUCC($k=8$) with strong symmetric protocol clearly demonstrates its superiority over ($k=4$ and $k=6$) as well as over weak symmetric ($k=8$) protocol, in predicting more accurate state and energy, as it achieves a fidelity very close to one in less than 200 function evaluations for both molecules. Thus, in the subsequent discussions, we use the NI-DUCC wavefunction, at $k=8$, employing the fourth protocol outlined earlier, which combines the strong symmetric excitations with a closure Lie algebraic structure.
@@ III.2 Energy dissociation curves and Energy convergence
Figures 3, 3, and 3 show the energy dissociation curves and the absolute values for the ground-state energy estimates, for LiH, H6, and BeH2, respectively, obtained by using NI-DUCC-VQE, at $k=8$. The same figures also include the energies of the FCI and untrotterized UCCSD-VQE methods.
The three methods yield similar energy estimates, with slight differences observed between UCCSD-VQE and NI-DUCC-VQE at the dissociation bond lengths of H6, and BeH2 molecules.
To make the differences between these methods more evident, figures 3, 3, and 3 present the differences with the exact FCI energy, highlighting the errors of UCCSD and NI-DUCC methods. Additionally, to perform a meaningful comparison, the energy convergence plots comparing these two methods with the unitary selective coupled cluster terminated at quadruple excitations (UsCCSDTQ) [31] are provided. UsCCSDTQ-VQE computations were terminated when reaching an iterative threshold $\epsilon=10^{-8}$ Hartree. We also included UsCCD-VQE energies using the same threshold. The UsCCD method generates the “strongly” double fermionic excitation orbitals, which is used in the first step while constructing the NI-DUCC-VQE algorithm. The UsCCD-VQE achieves chemical accuracy only over the equilibrium configurations, for the three molecules. UCCSD-VQE achieves chemical accuracy over all bond distances for LiH (3), and at bond distances close to the equilibrium configuration for H6 and BeH2. However, UCCSD-VQE struggles to achieve chemical accuracy for H6 and BeH2 when bond distances deviate from their equilibrium configurations, i.e. when the ground states become more strongly correlated.
@@ III.3 Accuracy Limits: NI-DUCC-VQE versus UsCCSDTQ-VQE and Trotterized sym-UCCSDT-VQE
The UsCCSDTQ method [31] is designed to recover additional correlation energy by including triple and quadruple excitations. Gradually adding these excitations has been shown to further improve the energy beyond the UCCSD method. However, we observed that, applying UsCCSDTQ-VQE to H6 and BeH2 molecules has some limitations in improving accuracy towards the reference FCI energy solution. This could be due to the crude approximation used for the screening coefficient computation [31]: $|H_{\beta}.t^{*}_{\beta}|\geq\epsilon$. Such an approximation is similar to the condition discussed earlier in Section I.2, where we first select the dominant fermionic double excitations in NI-DUCC-VQE. As a result, some important excitations appear to be missed during the iteration process, and thus the unitary selective approach, even at the quadruple excitation level, fails to improve the overall accuracy.
@@
In reference [31], UsCCSDTQ-VQE was tested on the BeH2 molecule at $r=3.9$( see Fig. 2.C) and it was observed that there is a region where the error remains constant (i.e without improvement in terms of chemical accuracy), despite the addition of more operators to the UsCCSDTQ ansatz. This indicates that some unimportant operators are being added in later iterations ahead of those that would have been more important, to reduce the energy error. This highlights the limitations of the UsCCDTQ approach. Therefore, increasing the iterative threshold ($\epsilon$) further to lower the overall UsCCDTQ-VQE energy appears no necessary.
@@
Similarly, the BeH2 molecule was simulated using the Trotterized sym-UCCSDT method [30] which exploits both spin and point group symmetries. According to our simulations shown in Fig. 1.B in [30], the sym-UCCSDT-VQE deviates by approximately 10-3 Hartree, and does not show further improvements. We suggested in the article that the sym-UCCSDT approach might miss important higher-order excitations, i.e beyond triple, which could be actually contributing significantly to the dynamical correlation, and therefore their absence leads to errors. A similar behavior was observed when sym-UCCSDTQ-VQE was tested on the H2O molecule (see Table 3 in [30]). Indeed these benchmarks demonstrate the challenge of studying highly correlated systems, where many electronic configurations significantly contribute to the wavefunction.
@@
In general, the limitations of triple and quadruple excitations at large bond lengths, could be potentially addressed by adding higher-order excitations, such as quintuple, sextuple, or septuple. However, this would require substantial effort in VQE optimization, due to the increased number of parameters and to resulting longer circuit depth, which is undesirable for current NISQ devices. Overall, UsCCSDTQ-VQE and sym-UCCSDT-VQE can construct ansatze to accurately approximate strongly correlated states, but, they do not actually guarantee similar high-level accuracy to the NI-DUCC ansatz. Overall, NI-DUCC-VQE demonstrates significant strength in covering strong correlations at large bond lengths, with high level of chemical accuracy, deviating by approximately 10-10 Hartree from FCI solution, as illustrated in the Figures above.
@@ IV Computational costs of the NI-DUCC-VQE algorithm
IV.1 NI-DUCC-VQE versus ADAPT-VQE

### v3

@@ II Computational Details
To generate the symmetry-preserving MCP excitations, which is the final step of the fourth component needed for the completion of the NI-DUCC algorithm, we used an additional python code, that generates an MCP by ensuring that the pool of excitations meets the required completeness conditions. This involves computing the product group and building their corresponding Lie algebra (see the python code given in [65]). It is important to note that computing the product group of a given pool is resource-intensive, especially as the number of qubits increases. Consequently, Section IV.3 provides detailed information about the algorithm computational resource demands while increasing the number of qubits. Finally, to address this issue in order to better estimate the computational requirements, we converted completely the python code into a more efficient C++ capable of a faster MCP generation.
@@ III.1 Numerical Simulations
Figure 2: Energy convergence plots for ground states of H6 (12 qubits) and BeH2 (14 qubits), using the STO-3G basis set, at bond distances ( $r_{\text{H-H}}$ = 1.0 Å,) and ($r_{\text{Be-H}}$ = 3.5 Å), respectively. The plots are obtained using the NI-DUCC-VQE algorithm: Top panel (a, b), compares the quality of excitation set, insuring the significance of strong symmetric excitations when they are combined with the closure Lie algebraic properties. In Figure 2(a), the weak symmetric excitations with or without the Lie algebraic closure are overlapping. Panel (c, d), illustrates the rapid convergence as the number of layers, $k$, increases in the circuit. Bottom Panel (e, f), shows the fidelity property, which is calculated by overlapping $\langle\Psi_{j}(\vec{\theta^{*}})|\Psi_{g}\rangle$ between the computed NI-DUCC state at each optimization step “$j$” and the theoretical eigenvector ground state $|\Psi_{g}\rangle$ of Hamiltonian ($\hat{H}$).
@@
In the last protocol, where strongly symmetric excitations are considered as starters and incorporate Lie closure relations, we observe that NI-DUCC-VQE allows for a “full” convergence of the H6 and BeH2 ground states, with an error of less than 10-10 Hartree relative to the reference FCI energy. The convergence is very rapid, requiring only about 800 function evaluations. These observations suggest that strong symmetric excitations with closure algebraic relations, might approximate strongly correlated states, much better than the NI-DUCC ansatz formed from strongly symmetric excitations missing Lie algebraic closure relations. Moreover, since we simulated BeH2 at a larger bond length of 3.5 Å, where the ground state is more strongly correlated, we expected to see some difference in the convergence behavior with respect to H6, which was tested near equilibrium at r= 1.0 Å. However, this was not the case, indicating that the Lie algebraic closure relations help to overcome this challenge. This demonstrates that incorporating Lie algebraic closure with strong symmetric starters in the NI-DUCC wavefunction systematically provides a solution to the ordering problem. Indeed, several studies [29, 70, 71] demonstrate that addressing ordering problems can enhance convergence and improve UCC-VQE energy. Similarly, we have shown improvements in the NI-DUCC accuracy thanks to the use of Lie algebraic closure relations.
We then further investigated the importance of using the “fourth” protocol in constructing the NI-DUCC wavefunction at $k=8$. As depicted in Figure 2 for H6 and Figure 2 for BeH2, we included energy convergence plots not only for $k=8$, but also for $k=3$ to $k=7$. We observed that when $k=3$ and $k=4$ are used, NI-DUCC approaches chemical accuracy (10-3 Hartree), while when using $k=5$, $k=6$ and $k=7$, it achieves an accuracy about $10^{-5}$ Hartree. Although with $k=8$, NI-DUCC-VQE approaches exact FCI solution (with an error less than 10-11 Hartree) after approximately 900 function evaluations without revealing any long plateaus, the differences between various $k$ layers may only matter, if one stops the calculation before a full convergence is reached. This could be the case, for instance, if achieving chemical accuracy is sufficient for a particular molecule. However, if the aim of the computation is to reach full convergence with the highest level of accuracy, then, using NI-DUCC excitations with sufficient number of layers is advantageous. Additionally, we have plotted (see Figure. 2 for H6 and Figure. 2 for BeH2), the overlap (“fidelity”) of the NI-DUCC-VQE state at each optimization step with the full configuration interaction (FCI) state, against the number of function evaluations (optimization steps). Various NI-DUCC tests were performed using different setups: (i) when strong symmetric starters are within closure Lie algebraic relations, we computed NI-DUCC($k=4$), NI-DUCC($k=6$), NI-DUCC($k=8$); (ii) when weak symmetric starters are within closure Lie algebraic relations, we computed NI-DUCC($k=8$). We observed that the NI-DUCC($k=8$) with strong symmetric protocol clearly demonstrates its superiority over ($k=4$ and $k=6$) as well as over weak symmetric ($k=8$) protocol, in predicting more accurate state and energy, as it achieves a fidelity very close to one in less than 200 function evaluations for both molecules. Thus, in the subsequent discussions, we use the NI-DUCC wavefunction, at $k=8$, employing the fourth protocol outlined earlier, which combines the strong symmetric excitations with a closure Lie algebraic structure.
@@ III.2 Energy dissociation curves and Energy convergence
Figures 3, 3, and 3 show the energy dissociation curves and the absolute values for the ground-state energy estimates, for LiH, H6, and BeH2, respectively, obtained by using NI-DUCC-VQE, at $k=8$. The same figures also include the energies of the FCI and untrotterized UCCSD-VQE methods. The three methods yield similar energy estimates, with slight differences observed between UCCSD-VQE and NI-DUCC-VQE at the dissociation bond lengths of H6, and BeH2 molecules.
To make the differences between these methods more evident, figures 3, 3, and 3 present the differences with the exact FCI energy, highlighting the errors of UCCSD and NI-DUCC methods. Additionally, to perform a meaningful comparison, the energy convergence plots comparing these two methods with the unitary selective coupled cluster terminated at quadruple excitations (UsCCSDTQ) [31] are provided. UsCCSDTQ-VQE computations were terminated when reaching an iterative threshold $\epsilon=10^{-8}$ Hartree. We also included UsCCD-VQE energies using the same threshold. The UsCCD method generates the “strongly” double fermionic excitation orbitals, which is used in the first step while constructing the NI-DUCC-VQE algorithm. The UsCCD-VQE achieves chemical accuracy only over the equilibrium configurations of the three molecules. Alternatively, UCCSD-VQE achieves chemical accuracy over all bond distances for LiH (3), and at bond distances close to the equilibrium configuration for H6 and BeH2. However, UCCSD-VQE struggles to achieve chemical accuracy for H6 and BeH2 when bond distances exceed 1.5 Å, where the ground states become more strongly correlated.
@@ III.3 Accuracy Limits: NI-DUCC-VQE versus UsCCSDTQ-VQE and Trotterized sym-UCCSDT-VQE
The UsCCSDTQ method [31] is designed to recover additional correlation energy by including triple and quadruple excitations. Gradually adding these excitations has been shown to further improve the energy beyond the UCCSD method. However, we observed that, applying UsCCSDTQ-VQE to H6 and BeH2 molecules has some limitations in improving accuracy towards the reference FCI energy solution.
This could be due to the crude approximation used for the screening coefficient computation [31]: $|H_{\beta}.t^{*}_{\beta}|\geq\epsilon$. Such an approximation is similar to the condition discussed earlier in Section I.2, where we first select the dominant fermionic double excitations in NI-DUCC-VQE. As a result, some important excitations appear to be missed during the iteration process, and thus the unitary selective approach, even at the quadruple excitation level, fails to improve the overall accuracy.
In reference [31], UsCCSDTQ-VQE was tested on the H6 molecule at bond distances $r=0.9$, $1.8$, and $3.6$ (see Fig. 2a). It was observed that accuracy challenges worsened as $r$ increased, with UsCCSDTQ deviating further from the FCI solution at distances beyond the equilibrium ($r=0.9$). In contrast, NI-DUCC was able to reach the FCI solution consistently across all tested distances.
@@
In the same reference [31], UsCCSDTQ-VQE was tested on the BeH2 molecule at $r=3.9$( see Fig. 2.C) and it was observed that there is a region where the error remains constant (i.e without improvement in terms of chemical accuracy), despite the addition of more operators to the UsCCSDTQ ansatz. This indicates that some unimportant operators are being added in later iterations ahead of those that would have been more important, to reduce the energy error. This highlights the limitations of the UsCCDTQ approach. Therefore, increasing the iterative threshold ($\epsilon$) further to lower the overall UsCCDTQ-VQE energy appears no necessary.
@@
Similarly, the BeH2 molecule was simulated using the Trotterized sym-UCCSDT method [30] which exploits both spin and point group symmetries. According to our simulations shown in Figure 1.B of reference [30], the sym-UCCSDT-VQE deviates by approximately 10-3 Hartree, and does not show further improvements. We suggested in the article that the sym-UCCSDT approach might miss important higher-order excitations, i.e beyond triple, which could be actually contributing significantly to the dynamical correlation, and therefore their absence leads to errors. A similar behavior was observed when sym-UCCSDTQ-VQE was tested on the H2O molecule (see Table 3 in [30]). Indeed these benchmarks demonstrate the challenge of studying highly correlated systems, where many electronic configurations significantly contribute to the wavefunction.
@@
In general, the limitations of triple and quadruple excitations at large bond lengths, could be potentially addressed by adding higher-order excitations, such as quintuple, sextuple, or septuple. However, this would require substantial effort in the VQE classical optimization, due to the increased number of parameters and to the resulting longer circuit depth, which is undesirable for current NISQ devices. Overall, UsCCSDTQ-VQE and sym-UCCSDT-VQE can construct ansätze to accurately approximate strongly correlated states, but, they do not actually guarantee similar high-level accuracy to the NI-DUCC ansatz. As discussed in Section III.1, the Lie algebraic closure with robust starter excitations in the NI-DUCC wavefunction, effectively addresses the ordering problem, in particular at extended bond lengths in H6 and BeH2, thus enhancing the NI-DUCC methodology relative to other ansätze (UsCCSDTQ and sym-UCCSDT) which do not guarantee ordering operators, and thus lead to convergence issues. Overall, NI-DUCC-VQE shows substantial strength in capturing strong correlations at large bond lengths, achieving chemical accuracy within approximately 10-10 Hartree of the FCI solution, as illustrated in the figures above.
@@ IV Computational costs of the NI-DUCC-VQE algorithm
IV.1 NI-DUCC-VQE versus fixed ansätze: sym-UCCSDT, UsCCSDTQ, COMPASS and COMPACT
After establishing NI-DUCC’s high accuracy compared to ansätze with higher-order fermionic excitations (up to quadrupole), we now turn to a detailed comparison. In this subsection, we evaluate NI-DUCC’s parameter efficiency and CNOT gate usage across different molecular systems, contrasting it with UsCCSDTQ and sym-UCCSDT (previously defined in the text), as well as two emerging fixed ansatz methods, COMPASS[48] and COMPACT [49]. Below, we provide a brief definition of the COMPASS and COMPACT methods:
  1. i.
    The COMPASS ansatz is a dynamically disentangled unitary coupled-cluster protocol that constructs an efficient ansatz using one- and two-body cluster operators alongside selected rank-two scatterers $\sigma$. Its accuracy is tunable through two threshold parameters, ($\epsilon_{1},\epsilon_{2}$), that control the inclusion of relevant operators. The ansatz has two versions based on the choice of the $\sigma$ operator in the selection bath: (a) the opposite-spin (OP) sector of $\sigma$ and (b) the partially paired (PP) sector, termed COMPASS-OP and COMPASS-PP, respectively (see Section III in [48]). Both versions can utilize parallel quantum architectures with energy-based sorting and commutativity prescreening to enhance efficiency and minimize gate depth.
  2. ii.
    COMPACT ansatz is a threshold-based, noise-robust strategy built on ab-initio many-body perturbation theory, and does not require any pre-circuit measurements. Its accuracy and quantum complexity are tunable by setting a perturbative order and threshold parameters, ($\epsilon_{1},\epsilon_{2},\epsilon_{3}$), which filter for the most significant excitations. This structure decomposes high-rank excitations into low-rank operators, which helps in optimizing computational resources and in minimizing gate depth.
We tested the efficacy and accuracy of the NI-DUCC method on three prototypical systems: LiH (at $r=1.0\,\text{\AA}$), BeH2 (at $r=2.0\,\text{\AA}$), and BH (at $r=2.0\,\text{\AA}$). These molecules and bond lengths were selected to enable a fair comparison with the COMPASS[48] and COMPACT [49] ansätze by drawing on data from their respective Figure 2.
As shown in Table 1, UsCCSDTQ achieves high accuracy ($10^{-7}$) with only 61 parameters, but at the cost of a very high CNOT gate count (26,608). In contrast, the sym-UCCSDT ansatz maintains accuracy with only 58 parameters and significantly fewer CNOT gates (14,304). Notably, COMPACT(5,5,4) requires just 44 parameters and achieves a much lower gate count (3,580) compared to sym-UCCSDT, underscoring its efficiency. The NI-DUCC method exhibits a progressive improvement in chemical accuracy with increasing layer depth $k$: from $10^{-5}$ at $k=6$, to $10^{-6}$ at $k=7$, and to $10^{-12}$ at $k=8$. Additionally, NI-DUCC achieves a significant reduction in gate count approximately by factor 3 compared to COMPACT(5,5,4) although the parameter count increases with additional layers.
For BeH2 (See Table 2), the sym-UCCSDT ansatz achieves a chemical accuracy of only $10^{-4}$ relative to FCI, while UsCCSDTQ improves this by two orders of magnitude, highlighting the impact of quadruple excitations. However, higher-order excitations also lead to a significant increase in the CNOT gate count as noticed in Table 2. In contrast, COMPASS-OP(5,5,4) and COMPASS-PP(5,5,4), which each uses up to two-body cluster operators, require only 54 parameters and reduce the CNOT gate count by approximately an order of magnitude compared to sym-UCCSDT. The NI-DUCC ansatz, at $k=6$ and $k=8$, improves accuracy from $10^{-5}$ to $10^{-7}$ with a minimal number of CNOT gates. At $k=8$, it reaches an accuracy of $10^{-12}$ with only 1,880 CNOT gates, in contrast to COMPASS-(5,5,4), which requires 2,592 gates to achieve $10^{-5}$, where 7 orders of magnitudes differ. However, NI-DUCC at this level necessitates roughly five times more parameters than COMPASS-(5,5,4).
The results for BH (Table 3) also demonstrate the NI-DUCC ansatz’s efficiency in minimizing CNOT gate counts compared to COMPASS and COMPACT. For an accuracy of $10^{-4}$, COMPASS-(5,7)-OP requires 2,736 CNOT gates, whereas NI-DUCC at $k=4$ achieves the same accuracy with only 680 CNOTs. Increasing NI-DUCC layers to $k=5$ improves accuracy by an order of magnitude with just 850 CNOTs. For $10^{-6}$ Hartree, COMPACT(3,3,4) requires 5,048 CNOTs and 59 parameters, while NI-DUCC at $k=6$ achieves this accuracy with 1,020 CNOTs and 204 parameters. NI-DUCC further enhances accuracy to $10^{-12}$ with only 1,360 CNOTs, considerably fewer than COMPASS or COMPACT, though it demands a higher parameter count. These results highlight NI-DUCC’s strong performance in reducing CNOT gates while maintaining a high level of chemical accuracy in the simulated molecules, outperforming the fixed ansatz approaches. However, they also indicate a need to further optimize the parameter count, which could potentially be improved with future techniques.
Molecule UsCCSDTQ sym-UCCSDT COMPACT(5,5,4) NI-DUCC($6$) NI-DUCC($7$) NI-DUCC($8$)
LiH (61) (26608) $10^{-7}$ (58)(14304) $10^{-7}$ (44) (3580) $10^{-7}$ (138) (690) $10^{-5}$ (161) (805) $10^{-6}$ (184) (920) $10^{-12}$
Table 1: Comparison of sym-UCCSDT, UsCCSDTQ, COMPACT(5,5,4) and NI-DUCC at layers $k=6,7,8$ for BeH2 geometry at bond length $r_{\text{Li-H}}$ = 1.0 Å in the STO-3G basis. Estimations provided in each box include parameter count, CNOT count, and energy error relative to FCI (Hartee) (Same for Table II and Table III). Data for UCCSDT, sym-UCCSDT and COMPACT(5,5,4) are extracted from [49], where COMPACT uses thresholds $\log(\varepsilon_{1})$, $\log(\varepsilon_{2}$) and $\log(\varepsilon_{3}$) to select significant spin-complementary excitations in the ansatz. Our UsCCSDTQ simulations are terminated at a threshold of $\epsilon=10^{-8}$ Hartree. CNOT count scaling details are available in Supplementary Table 1 of Supplementary Materials [59] for all ansätze.
Molecule UsCCSDTQ sym-UCCSDT COMPASS(5,7)-OP COMPASS(5,7)-PP NI-DUCC($6$) NI-DUCC($7$) NI-DUCC($8$)
BeH2 (84) (44840) $10^{-6}$ (146)(21696) $10^{-4}$ (54) (2592) $10^{-5}$ (54) (2592) $10^{-5}$ (282) (1410) $10^{-5}$ (329) (1645) $10^{-6}$ (376) (1880) $10^{-12}$
Table 2: Comparison of UsCCSDTQ, COMPASS(5,7)-OP, COMPASS(5,7)-PP, and NI-DUCC at layers $k=6,7,8$ for BeH2 geometry at bond length $r_{\text{Be-H}}$ = 2.0 Å in the STO-3G basis. Data for COMPASS(5,7)-OP and COMPASS(5,7)-PP, are extracted from [48], where COMPASS uses thresholds $\log(\varepsilon_{1})$ and $\log(\varepsilon_{2}$) to select significant cluster amplitudes and scatterers in the ansatz.
Molecule COMPASS(5,7)-OP COMPACT(3,3,4) NI-DUCC($4$) NI-DUCC($5$) NI-DUCC($6$) NI-DUCC($7$) NI-DUCC($8$)
BH (57) (2736) $10^{-4}$ (59) (5048) $10^{-6}$ (136) (680) $10^{-4}$ (170) (850) $10^{-5}$ (204) (1020) $10^{-6}$ (238) (1190)$10^{-7}$ (272) (1360) $10^{-12}$
Table 3: Comparison of COMPASS(5,7)-OP COMPACT(3,3,4) and NI-DUCC at layers $k=4,5,6,7,8$ for BH geometry at bond length $r_{\text{B-H}}$ = 2.0 Å in the STO-3G basis. Data for COMPASS(5,7)-OP and COMPACT(3,3,4) are extracted from figure 2 of [48] and [49] respectively.
IV.2 NI-DUCC-VQE versus ADAPT-VQE